Estimate the stiffness (k) of a spring in a child’s pogo stick if the child has a mass of 35 kg and bounces once every 2.0 seconds. [ANS = 350 N m-1]


1
Expert's answer
2021-11-25T16:23:08-0500

Given:

"m=35\:\rm kg"

"T=2.0\:\rm s"


The period of oscillation of a spring pendulum

"T=2\pi \sqrt{\frac{m}{k}}"

Hence, the stiffness of a spring

"k=\frac{4\pi^2m}{T^2}=\frac{4*3.14^2*35}{2.0^2}=345.4\:\rm N/m\approx 350 \:N/m"